# Energy Weighted Learning Progress Guided Interleaved Multi-Task Learning

### v1

@@ III Problem Statement
The problem of interest is to answer how a dynamical task arbitration mechanism for data sampling and learning can lead to better overall learning and skill transfer in multitask learning. The multitask learning framework we consider follows human learning, where the learner has to decide which task to engage in and when to disengage, effectively allowing the emergence of arbitrary interleaved learning regimes.
Specifically, in the context of supervised regression problems, we design an online multi-task learning model equipped with a task arbitration mechanism that enables the agent to decide, at each training step, which task to focus on for data sampling and learning. Formally, given $T$ tasks where task $t\in\{1,2,\dots,T\}$, in each training step, the task arbitration mechanism chooses a task $t^{*}\in T$ to be learned by the multi-task learning agent. Next, the agent interacts with the corresponding task environment $\mathcal{E}_{t^{*}}$ and collects new data
$\mathcal{D}_{t^{*}}=\{(x_{i}^{(t^{*})},y_{i}^{(t^{*})})\}_{i=1}^{N_{t^{*}}},$
where $x_{i}^{(t^{*})}$ represents the $i$-th input for task $t^{*}$, $y_{i}^{(t^{*})}$ is the corresponding target and $N_{t^{*}}$ is the number of samples of task $t^{*}$. This approach contrasts with the standard offline setup, where a complete dataset is available prior to the beginning of training. Once the agent gathers the $\mathcal{D}_{t^{*}}$, it then updates its specific parameter set $\theta^{(t^{*})}$ along with the $\theta^{s}$ which is the parameter set shared among the tasks. This dynamic process continues until the end of the training procedure. The predictive function for each task $t\in T$ is given by
$f_{t}\bigl{(}x^{(t)};\theta^{s},\theta^{(t)}\bigr{)}.$
Each task is associated with a loss function $\mathcal{L}_{t}$, which quantifies the error between the predicted outputs and the true labels. The objective is to find parameters $\theta^{s},\{\theta^{(t)}\}_{t=1}^{T}$ that minimize the total loss across all tasks:
$\min_{\theta^{s},\{\theta^{(t)}\}}\sum_{t=1}^{T}\mathcal{L}_{t}\left(\left\{\left(f_{t}(x_{i}^{(t)};\theta^{s},\theta^{(t)}),\,y_{i}^{(t)}\right)\right\}_{i=1}^{N_{t}}\right).$
@@
Coming to the effect prediction tasks in this work, an agent is required to learn a total of $T$ distinct effect prediction tasks by establishing a mapping from the state-action space to the effect space. For each task $t\in\{1,2,\ldots,T\}$, the state is represented as $x^{(t)}\in\mathbb{R}^{n_{t}}$ and the action as $a^{(t)}\in\mathbb{R}^{m_{t}}$, where $n_{t}$ and $m_{t}$ denote the dimensions of the state and action spaces, respectively. The corresponding effect, which is the outcome of executing action $a^{(t)}$ in state $x^{(t)}$, is denoted by $e^{(t)}\in\mathbb{R}^{k_{t}}$. The goal is to dynamically arbitrate data sampling and learning among the tasks so that for each task $t$ action-effect prediction functions, $f_{t}:\mathbb{R}^{n_{t}}\times\mathbb{R}^{m_{t}}\rightarrow\mathbb{R}^{k_{t}}$, are learned. After learning, it is expected that for any state-action pair $(x^{(t)},a^{(t)})$ the predicted effect $\tilde{e}^{(t)}=f_{t}(x^{(t)},a^{(t)})$ closely approximates the actual effect $e^{(t)}$. To formalize this further, we can define the data set an agent is able to collect with
$\mathcal{D}_{t}=\{(x_{i}^{(t)},a_{i}^{(t)},e_{i}^{(t)})\}_{i=1}^{N_{t}}$
where each tuple provides a sample of the state, action, and resulting effect.
IV Methodology
Figure 1: Overview of the proposed multi-task effect prediction architecture and an example training iteration for $t^{*}=2$ where total number of tasks $T=3$. The model features shared and task-specific components, with only the selected task being trained at each step based on the arbitration mechanism.
In this section, we detail the architecture and methodologies employed in our proposed interleaved multi-task learning framework with energy-modulated learning progress. Our goal is to realize human-like learning by interleaving tasks and optimizing energy consumption through a specialized neural network architecture and task selection mechanism.
IV-A Neural Network Architecture
@@
Our model is based on an encoder-decoder network architecture designed to handle multiple-task learning while allowing for inter-task information sharing. The key components of the architecture are task-specific input (state) and action projection layers, task-specific encoders as well as a shared encoder and a shared attention module, and finally task specific output (effect) decoders. We detail these components next.
@@
    State Projection Layer. For each task $t$ in the set of tasks $T$, the state input $x^{(t)}\in\mathbb{R}^{n_{t}}$ is first passed through a task-specific projection function $P_{state}^{(t)}\colon\mathbb{R}^{n_{t}}\to\mathbb{R}^{d_{s}}$ where $n_{t}$ is the input dimensionality of task $t$ and $d_{s}$ is the dimensionality of the projected state input. This layer maps the input into a fixed-dimensional space to ensure consistency across different tasks. Mathematically, the projected input is given by:
    $x^{\prime(t)}=P_{state}^{(t)}(x^{(t)}),$ (1)
    where $x^{\prime(t)}\in\mathbb{R}^{d_{s}}$. This step ensures that inputs from different tasks are compatible for subsequent shared processing. For the implementation of this layer, either a linear or nonlinear function, depending on the specific problem, can be chosen. However, since the primary objective of these layers is to map inputs of varying dimensions onto a uniform dimensional space, a linear projection function is implemented in our work, as often done in the literature [87, 88].
@@
    Action Projection Layer. Since the action vectors $a^{(t)}\in\mathbb{R}^{m_{t}}$ for each task $t$ can vary in dimensionality, similar to the state projection layers, we project them to a fixed-dimensional space to maintain consistency across tasks. Each action vector is passed through a task-specific projection function $P_{action}^{(t)}\colon\mathbb{R}^{m_{t}}\to\mathbb{R}^{d_{a}}$:
    $a^{\prime(t)}=P_{action}^{(t)}(a^{(t)}),$ (2)
    where $a^{\prime(t)}\in\mathbb{R}^{d_{a}}$ is the projected action vector for task $t$ and $d_{a}$ is the fixed dimensionality of the projected action space. This projection ensures that actions from different tasks are compatible for further processing. Similar to the (1), we choose to have a linear projection layer for the action vectors.
@@
    Shared State Encoder. We incorporate a shared encoder in our neural network with the motivation that it will extract general features that are beneficial across multiple tasks, thus promoting knowledge sharing while reducing redundancy, leading to more efficient learning. The projected input $x^{\prime(t)}$ is then fed into the a shared encoder $F\colon\mathbb{R}^{d_{s}}\to\mathbb{R}^{d_{h}}$, to extract low-level features common across different tasks:
    $h=F(x^{\prime(t)}),$ (3)
    where $h\in\mathbb{R}^{d_{h}}$ is the encoded representation, and $d_{h}$ is the dimensionality of the shared encoder output.
@@
    Task State Encoders. Following the shared encoder, further, we integrate task-specific encoders aiming to capture state knowledge unique to each task that the shared encoder cannot extract. We hypothesize that each task benefits from specialized features, enhancing its performance during training. Hence, the $h$ is forwarded through all task-specific encoders in order to generate task-specific latent representations:
    $r^{(t)}=f^{(t)}(h),\quad\forall t\in T$ (4)
    where $r^{(t)}\in\mathbb{R}^{d_{r}}$ is the task-specific latent representation, $f^{(t)}\colon\mathbb{R}^{d_{h}}\to\mathbb{R}^{d_{r}}$ is the task-specific encoder function of $t$ and $d_{r}$ is the dimensionality of the task specific latent space. Note that during the training of task $t^{*}$, for all $f^{(t)}$ where $t\neq t^{*}$ are frozen, which means gradient propagation is not carried out for them.
@@
    Shared Attention Module. To facilitate inter-task communication and let the network focus on the current training task, we utilize a multi-head attention (MHA) mechanism [89]. Instead of having one attention function, where $Q,K,V\in\mathbb{R}^{d_{model}}$, MHA layer computes attention across $H$ parallel heads that each head projects $Q,K,V$ into a lower-dimensional space and performs scaled dot-product attention. Let $d_{k}$ be the dimensionality per head, and let $W_{i}^{Q},W_{i}^{K},W_{i}^{V}\in\mathbb{R}^{d_{model}\times d_{k}}$ be the projection matrices for head $i\in\{1,\dots,H\}$. Then, following [89] we have:
@@
    $A=\left[\mathrm{head}_{1}\;:\;\mathrm{head}_{2}\;:\;\dots\;:\;\mathrm{head}_{H}\right]W^{O}\in d_{model},$
@@
    • Beforehand the attention function, we concatenate an additional flag bit to each task-specific latent representation, $r^{(t)}$. Specifically, for each task $t$, we concatenate a training flag $\delta^{(t)}$ indicating whether task $t$ is currently being trained $(\delta^{(t)}=1)$ or not $(\delta^{(t)}=0)$. The concatenated vector for each task is:
      $z^{(t)}=[r^{(t)}\;:\;\delta^{(t)}]\in\mathbb{R}^{d_{r}+1}$ (5)
      From now on, the $z^{(t)}$ here is the final representation of task $t$.
    • After gathering the $z^{(t)}$ for each task, they are stacked to form a matrix $Z$:
      $Z=\begin{bmatrix}z^{(1)}\\ z^{(2)}\\ \vdots\\ z^{(T)}\end{bmatrix}\in\mathbb{R}^{T\times(d_{r}+1)}$ (6)
      where $T$ is the number of tasks, $d_{r}$ is the fixed dimensionality of task-specific latent representations. Extra dimension accounts for the training flag.
    • Once we obtain the matrix $Z$, the $Q,K$ and $V$ values are gathered as below:
@@
      $Q=z^{(t^{*})}\in\mathbb{R}^{(d_{r}+1)},$
      where $d_{r}+1$ is the embedding dimension, or $d_{model}$.
@@
      $K=V=Z\in\mathbb{R}^{T\times(d_{r}+1)},$
    Using these $Q,K,V$ values, we feed them into our shared MHA module and get the attention output $A$ for the task $t=t^{*}$:
    $A^{(t)}=SharedMHA(Q,K,V),$ (7)
    where $A^{(t)}\in\mathbb{R}^{(d_{r}+1)}$. Because the query is restricted to the representation of the active task, $z^{(t^{*})}$, the network focuses on how the other tasks’ representations (contained in $Z$) can inform task $t^{*}$. Moreover, the flags within each row of $Z$ indicate whether a given input originates from the actively trained task ($\delta^{t=t^{*}}=1$) or from another task ($\delta^{t\neq t^{*}}=0$), further guiding the attention toward balancing shared and task-specific information.
@@
    Task Effect Decoders. Same as the task-specific sub-encoder $f^{(t)}$, for each task $t$; there is a decoder module $g^{(t)}$ designed for predicting the resulted output (effect) $e^{(t)}$, after the executed action $a^{(t)}$ while the environment is in the state $x^{(t)}$. For task $t=t^{*}$, the output from the shared MHA mechanism $A^{(t)}$ along with the projected action vector $a^{\prime(t)}$ from (2) is given to the task-specific decoder of the training task only, contrary to the (4), as follows:
    $\tilde{e}^{(t)}=g^{(t)}(A^{(t)},a^{\prime(t)}).$ (8)
    Here $\tilde{e}^{(t)}\in\mathbb{R}^{k_{t}}$ and $k_{t}$ is the output dimension of task $t$. By integrating contextual information from the shared attention mechanism with the projected action vector, the decoder maps the state information to the corresponding effect.
IV-B Task Arbitration
To investigate the effects of interleaving on multi-task learning, we propose a task selection strategy based on the dynamic assessment of each task’s learning progress and, additionally, its associated energy consumption. Unlike traditional multi-task learning approaches that train tasks in a fixed schedule or isolate them entirely, our method continuously evaluates the tasks’ performance trends and selects which task to train next based on these evaluations. By doing so, we simulate the human tendency to switch between tasks in response to learning progress and energetic constraints. Below, the interleaved task selection algorithm is explained in detail:

### v2

@@ 3 Problem Statement
The general problem addressed is how a task arbitration and skill transfer mechanism can lead to improved overall learning and skill transfer performance in multitask learning settings. The framework sought to mimic human learning, where the learner has to decide which task to engage in and when to disengage, effectively allowing the emergence of arbitrary interleaved learning regimes.
We focus on supervised learning problems in robotics, and seek an online multi-task learning model equipped with a task arbitration mechanism that enables the robot to decide which task to engage in for data sampling and learning with the goal of minimizing average prediction error over the tasks. Formally, we are given $m$ tasks $\left\{\mathcal{T}_{i}\right\}_{i=1}^{m}$ and an allocated interaction period $N_{\mathcal{T}_{i}}$ for each task. When task $\mathcal{T}_{i}$ is engaged, it is allowed to execute a total of $N_{\mathcal{T}_{i}}$ interactions with the environment. After this period, the task arbitration mechanism chooses a new task to be engaged. The selected task is used in execution, data collection, and learning. Denoting the selected task with $\mathcal{T}_{*}$, the execution experience for task $\mathcal{T}_{*}$ can be represented in the form of input, output pairs
$\mathcal{D}_{\mathcal{T}_{*}}=\{(x_{j}^{\mathcal{T}_{*}},y_{j}^{\mathcal{T}_{*}})\}_{j=1}^{N_{\mathcal{T}_{*}}},$
where $x_{j}^{\mathcal{T}_{*}}$ represents input while executing task $\mathcal{T}_{*}$, and $y_{j}^{\mathcal{T}_{i}}$ is the corresponding output at sampling step $j$. Note that by setting $N_{\mathcal{T}_{i}}$ values differently, one can encode the expected relative difficulty of the tasks, as in the case for tasks with high-dimensional versus low-dimensional input spaces. Once the agent gathers a mini-batch of experience, $\mathcal{D}_{\mathcal{T}_{*}}$, it updates the task specific parameters, $\theta^{\mathcal{T}_{*}}$ as well as the shared task parameters, $\theta^{s}$. This online sampling and learning process continues to minimize the prediction error for each task. Denoting the prediction for task $\mathcal{T}_{i}$ and loss associated with it by $f_{\mathcal{T}_{i}}\bigl{(}x^{\mathcal{T}_{i}};\theta^{s},\theta^{\mathcal{T}_{i}}\bigr{)}$, $\mathcal{L}_{\mathcal{T}_{i}}$ respectively, the objective for the multi-task learning agent can be stated as finding $\theta^{s},\{\theta^{\mathcal{T}_{i}}\}_{i=1}^{m}$ to minimize the total loss, $\mathcal{TL}$ across tasks:
$\mathcal{TL}=\min_{\theta^{s},\{\theta^{\mathcal{T}_{i}}\}}\sum_{i=1}^{m}\mathcal{L}_{\mathcal{T}_{i}}\left(\left\{\left(f_{\mathcal{T}_{i}}(x_{j}^{\mathcal{T}_{i}};\theta^{s},\theta^{\mathcal{T}_{i}}),\,y_{j}^{\mathcal{T}_{i}}\right)\right\}_{j=1}^{N_{\mathcal{T}_{i}}}\right).$
@@
The problem statement projected to the experiments conducted in this paper can be given as follows: The learning agent is required to learn a total of $m$ distinct effect prediction tasks by establishing a mapping from the state-action space to the effect space. For each task $\mathcal{T}_{i}$, the state is represented as $x^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{s}^{\mathcal{T}_{i}}}$ and the action as $a^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{a}^{\mathcal{T}_{i}}}$, where $d_{s}^{\mathcal{T}_{i}}$ and $d_{a}^{\mathcal{T}_{i}}$ denote the dimensions of the state and action spaces of task $\mathcal{T}_{i}$, respectively. The corresponding effect, which is the outcome of executing action $a^{\mathcal{T}_{i}}$ in state $x^{\mathcal{T}_{i}}$, is denoted by $e^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{e}^{\mathcal{T}_{i}}}$. Thus, the experience collected and used in learning is $\mathcal{D}_{\mathcal{T}_{i}}=\{((x_{j}^{\mathcal{T}_{i}},a_{j}^{\mathcal{T}_{i}}),e_{j}^{\mathcal{T}_{i}})\}_{j=1}^{N_{\mathcal{T}_{i}}}$. The goal is to dynamically arbitrate data sampling and learning among the tasks so that for each task $\mathcal{T}_{i}$ action-effect prediction functions, $f_{\mathcal{T}_{i}}:\mathbb{R}^{d_{s}^{\mathcal{T}_{i}}}\times\mathbb{R}^{d_{a}^{\mathcal{T}_{i}}}\rightarrow\mathbb{R}^{d_{e}^{\mathcal{T}_{i}}}$, are learned. After learning, it is expected that for any state-action pair $(x^{\mathcal{T}_{i}},a^{\mathcal{T}_{i}})$ the predicted effect $\tilde{e}^{\mathcal{T}_{i}}=f_{\mathcal{T}_{i}}(x^{\mathcal{T}_{i}},a^{\mathcal{T}_{i}})$ closely approximates the actual effect $e^{\mathcal{T}_{i}}$.
4 Methodology
In this section, we give the details of the learning architecture and task switching mechanisms developed and used in the experiments.
4.1 Neural Network Architecture
@@
Our model is based on an encoder-decoder network architecture designed to handle multiple-task learning while allowing for inter-task information sharing. The key components of the architecture are task-specific input (state) and action projection layers, task-specific encoders as well as a shared encoder and a shared attention module, and finally task-specific output (effect) decoders (see Figure 1). We detail these components next. Symbols and notations used are defined in Table 1 for clarity (see Table 1 in the Appendix for the actual sizes).
Figure 1: Overview of the proposed multi-task effect prediction architecture and an example training iteration for $\mathcal{T}_{*}$ where =2*=2∗ = 2 and total number of tasks $m=3$. The model features shared and task-specific components, with only the selected task being trained at each step based on the arbitration mechanism.
Symbol Description
$m$ Total number of tasks
$i$ Task index, $\{\mathcal{T}_{i}\}_{i=1}^{m}$
$\mathcal{T}_{*}$ Selected task to be trained
$x^{\mathcal{T}_{i}}$ State of task $\mathcal{T}_{i}$, $x^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{s}^{\mathcal{T}_{i}}}$
$a^{\mathcal{T}_{i}}$ Action of task $\mathcal{T}_{i}$, $a^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{a}^{\mathcal{T}_{i}}}$
$e^{\mathcal{T}_{i}}$ Effect of task $\mathcal{T}_{i}$, $e^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{e}^{\mathcal{T}_{i}}}$
$P^{\mathcal{T}_{i}}_{\text{state}}$ Task-specific state projection function, $P^{\mathcal{T}_{i}}_{\text{state}}:\mathbb{R}^{d_{s}^{\mathcal{T}_{i}}}\rightarrow\mathbb{R}^{d_{s}}$
$F$ Shared state encoder, $F:\mathbb{R}^{d_{s}}\rightarrow\mathbb{R}^{d_{h}}$
$h$ Output of the shared encoder, $h\in\mathbb{R}^{d_{h}}$
$r^{\mathcal{T}_{i}}$ Task-specific latent representation, $r^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{r}}$
$z^{\mathcal{T}_{i}}$ Task-specific latent representation with flag bit, $z^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{r}+1}$
$Z$ Stacked matrix of all $\{z^{\mathcal{T}_{i}}\}_{i=1}^{m}$, $Z\in\mathbb{R}^{m\times(d_{r}+1)}$
$P^{\mathcal{T}_{i}}_{\text{action}}$ Task-specific action projection function, $P^{\mathcal{T}_{i}}_{\text{action}}:\mathbb{R}^{d_{a}^{\mathcal{T}_{i}}}\rightarrow\mathbb{R}^{d_{a}}$
$f^{\mathcal{T}_{i}}$ Task-specific encoder for task $\mathcal{T}_{i}$, $f^{\mathcal{T}_{i}}:\mathbb{R}^{d_{h}}\rightarrow\mathbb{R}^{d_{r}}$
$g^{\mathcal{T}_{i}}$ Task-specific decoder for task $\mathcal{T}_{i}$, $g^{(t)}:\mathbb{R}^{d_{r}+d_{a}+1}\rightarrow\mathbb{R}^{d_{e}^{\mathcal{T}_{i}}}$
Table 1: Overview of symbols and notations.
@@
    State Projection Layer. For each task $\mathcal{T}_{i}$, the state input $x^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{s}^{\mathcal{T}_{i}}}$ is first passed through a task-specific projection function $P_{state}^{\mathcal{T}_{i}}\colon\mathbb{R}^{d_{s}^{\mathcal{T}_{i}}}\to\mathbb{R}^{d_{s}}$ leading to the projected input vector given by:
    $\hat{x}^{\mathcal{T}_{i}}=P_{state}^{\mathcal{T}_{i}}(x^{\mathcal{T}_{i}}),$ (1)
    where $\hat{x}^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{s}}$. This step ensures that inputs from different tasks have the same dimensionality for subsequent shared processing. The projection function may be chosen to be linear or nonlinear. As the main objective of these layers is to have the same dimension at the output, a linear projection is used in the current work, as often done in the literature [100, 101].
@@
    Action Projection Layer. Since the action vector $a^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{a}^{\mathcal{T}_{i}}}$ for each task $t$ can vary in dimensionality, similar to the state projection layers, we project them to a fixed-dimensional space to maintain consistency across tasks. Each action vector is passed through a task-specific projection function $P_{action}^{\mathcal{T}_{i}}\colon\mathbb{R}^{d_{a}^{\mathcal{T}_{i}}}\to\mathbb{R}^{d_{a}}$:
    $\hat{a}^{\mathcal{T}_{i}}=P_{action}^{\mathcal{T}_{i}}(a^{\mathcal{T}_{i}}),$ (2)
    where $\hat{a}^{\mathcal{T}_{i}}\in\mathbb{R}^{d_{a}}$ is the projected action vector for task $\mathcal{T}_{i}$. This projection ensures that actions from different tasks are compatible for further processing. Similar to the (1), we choose to have a linear projection layer for the action vectors.
@@
    Shared State Encoder. We incorporate a shared encoder in our neural network to encourage the extraction of general features that may be beneficial across multiple tasks, promoting knowledge sharing while reducing redundancy. Hence the projected input $\hat{x}^{\mathcal{T}_{i}}$ is fed into the a shared encoder $F\colon\mathbb{R}^{d_{s}}\to\mathbb{R}^{d_{h}}$, to extract a $d_{h}$-dimensional shared feature vector, $h\in\mathbb{R}^{d_{h}}$ with
    $h=F(\hat{x}^{\mathcal{T}_{i}}),$ (3)
@@
    Task State Encoders. Following the shared encoder, we further integrate task-specific encoders, aiming to capture state knowledge unique to each task that the shared encoder cannot extract. We hypothesize that each task benefits from specialized features, enhancing its performance during training. Hence, $h$ is forwarded through all task-specific encoders in order to generate task-specific latent representations:
    $r^{\mathcal{T}_{i}}=f^{\mathcal{T}_{i}}(h),\quad\{\mathcal{T}_{i}\}_{i=1}^{m}$ (4)
    Note that when training task $\mathcal{T}_{*}$, only the parameters of $f^{\mathcal{T}_{*}}$ are updated; all other $f^{\mathcal{T}_{i}}$ modules remain frozen, with no gradient propagation.
@@
    Shared Attention Module. To facilitate inter-task communication and let the network focus on the current training task, we utilize a multi-head attention (MHA) mechanism [102]. Instead of having one attention function, where $Q,K,V\in\mathbb{R}^{d_{h}}$, MHA layer computes attention across $H$ parallel heads that each head projects $Q,K,V$ into a lower-dimensional space and performs scaled dot-product attention. Let $d_{k}$ be the dimensionality per head, and let $W_{i}^{Q},W_{i}^{K},W_{i}^{V}\in\mathbb{R}^{d_{h}\times d_{k}}$ be the projection matrices for head $i\in\{1,\dots,H\}$. Then, $Q,K,V$ are projected as follows:
@@
    $A=\left[\mathrm{head}_{1}\;:\;\mathrm{head}_{2}\;:\;\dots\;:\;\mathrm{head}_{H}\right]W^{O}\in d_{h},$
@@
    • (a)
      Beforehand the attention function, we concatenate an additional flag bit to each task-specific latent representation, $r^{\mathcal{T}_{i}}$. Specifically, for each task $\mathcal{T}_{i}$, we concatenate a training flag $\delta^{\mathcal{T}_{i}}$ indicating whether task $\mathcal{T}_{i}$ is currently being trained $(\delta^{\mathcal{T}_{i}}=1)$ or not $(\delta^{\mathcal{T}_{i}}=0)$. The concatenated vector for each task is:
      $z^{\mathcal{T}_{i}}=[r^{\mathcal{T}_{i}}\;:\;\delta^{\mathcal{T}_{i}}]\in\mathbb{R}^{d_{r}+1}$ (5)
      From now on, we call $z^{\mathcal{T}_{i}}$ the final representation of task $\mathcal{T}_{i}$.
    • (b)
      After gathering $z^{\mathcal{T}_{i}}$ for each task, they are stacked to form a matrix $Z$:
      $Z=\begin{bmatrix}z^{\mathcal{T}_{1}}\\ z^{\mathcal{T}_{2}}\\ \vdots\\ z^{\mathcal{T}_{m}}\end{bmatrix}\in\mathbb{R}^{m\times(d_{r}+1)}.$ (6)
    • (c)
      Once we obtain the matrix $Z$; $Q,K$ and $V$ values are gathered as below:
@@
      $Q=z^{\mathcal{T}_{*}}\in\mathbb{R}^{(d_{r}+1)}.$
@@
      $K=V=Z\in\mathbb{R}^{m\times(d_{r}+1)}.$
    Using these $Q,K,V$ values, we feed them into our shared MHA module and get the attention output $A$ for the task $\mathcal{T}_{*}$:
    $A^{\mathcal{T}_{*}}=SharedMHA(Q,K,V),$ (7)
    where $A^{\mathcal{T}_{*}}\in\mathbb{R}^{(d_{r}+1)}$. Because the query is restricted to the representation of the active task, $z^{\mathcal{T}_{*}}$, the network focuses on how the other tasks’ representations (contained in $Z$) can inform task $\mathcal{T}_{*}$. Moreover, the flags within each row of $Z$ indicate whether a given input originates from the actively trained task ($\delta^{\mathcal{T}_{i=*}}=1$) or from another task ($\delta^{\mathcal{T}_{i\neq*}}=0$), further guiding the attention toward balancing shared and task-specific information.
@@
    Task Effect Decoders. Same as the task-specific sub-encoder $f^{\mathcal{T}_{i}}$, for each task $\mathcal{T}_{i}$; we utilize a decoder module $g^{\mathcal{T}_{i}}$ designed for predicting the resulting effect (output) $e^{\mathcal{T}_{i}}$, after the executed action $a^{\mathcal{T}_{i}}$ while the environment is in the state $x^{\mathcal{T}_{i}}$. For task $\mathcal{T}_{*}$, the output from the shared MHA mechanism $A^{\mathcal{T}_{*}}$ along with the projected action vector $\hat{a}^{\mathcal{T}_{*}}$ from (2) is given to the task-specific decoder of the selected task only, contrary to the (4), as follows:
    $\tilde{e}^{\mathcal{T}_{*}}=g^{\mathcal{T}_{*}}([A^{\mathcal{T}_{*}}:\hat{a}^{\mathcal{T}_{*}}]).$ (8)
    where $\tilde{e}^{\mathcal{T}_{*}}\in\mathbb{R}^{d_{e}^{\mathcal{T}_{*}}}$. By integrating contextual information from the shared attention mechanism with the projected action vector, the decoder maps the state information to the corresponding effect.
4.2 Task Arbitration
To investigate the effects of allowing interleaved learning in multi-task settings, we propose a task selection strategy based on the assessment of each task’s learning progress or its discounted version by the neural cost of learning. Unlike traditional multi-task learning approaches that train tasks in a fixed schedule or isolate them entirely, our method continuously evaluates the tasks’ performance trends and selects which task to train next based on these evaluations. By doing so, we aim to emulate the human tendency to switch among tasks. An overview of the proposed interleaved task engagement mechanism is presented in Algorithm 1, where a non-greedy selection strategy is adopted by giving other tasks a chance to be selected, with an exploration rate $\epsilon$, thereby reducing over-reliance on the highest-scoring task.